OpenStudy (anonymous):

Write 2 equations of a line that passes through (-3,-4) one is perpendicular and the other one is parallel to y=4x+1 show work

OpenStudy (anonymous):

Parallel: -4=4*(-3)+b ; b=-4+12; b = 8; then........... y=4x+8

OpenStudy (bradely):

step by step answers posted here http://www.mathskey.com/question2answer/1993/write-equations-line-that-passes-through-one-perpendicular?show=1994#a1994 ask your Algebra2 homework questions at http://www.mathskey.com/question2answer/algebra-2 and get free math help. all the best